Please help me find x and y.
3x-54=y
3y-54=0

Answer by stanbon(75887)

Please help me find x and y.
3x-54=y
3y-54=0
---
y = 18
---
Solve for "x":
3x-54 = 18
3x = 72
x = 24
=============
Cheers,
Stan H.